# A Survey of Text Classification Under Class Distribution Shift

### v2

@@ 4 Zero-Shot Text Classification
Formal definition. Zero-shot learning aims to recognize samples from $C^{uk}$ using knowledge transferred from $C^{kk}$. Formally, the training data is defined as $\mathcal{D}=\{(x,y)\mid x\in\mathcal{X},y\in C^{kk}\}$, while the test data is defined as set $\mathcal{T}=\{(x,y)\mid x\in\mathcal{X},y\in C^{uk}\}$. Knowledge is usually transferred via an auxiliary semantic space AA, e.g. word embeddings or attribute vectors, such that $f:\mathcal{X}\to A$ and $g:A\to\mathcal{Y}$. The mapping $f$ learns to project features into the semantic space, while $g$ maps semantic representations to class labels.
Trends and approaches. When no examples from the target classes are available, various alternative tasks can be leveraged, with or without additional training, to perform zero-shot text classification. For instance, Yin et al. (2019) propose using the entailment task as a zero-shot text classification task. To decide if a text $x$ should be classified into a certain class, e.g. “politics”, one can ask if $x$ entails “The previous text is about politics”. This procedure is applied for each of the possible classes, and the one with the best confidence is considered the correct one. Using this technique or similar ones, any model capable of solving the entailment problem can implicitly be used to perform zero-shot text classification.
Instruction tuning, defined as fine-tuning language models on a collection of datasets described via instructions, can be seen as a generalization over the framework proposed by Yin et al. (2019). Instruction tuning can lead to better performance on unseen tasks, if the target task is described via simple instructions provided in natural language. Models trained with instruction tuning can perform various tasks, including zero-shot classification, as shown by Zhang et al. (2024). This capability is achieved by training language models to respond to simple natural language commands, with some degree of generality. For instance, Wei et al. (2022) and Sanh et al. (2022) explore similar ideas, dividing the tasks for which datasets are available into separate clusters, each containing multiple datasets. Fine-tuning via instructions on any of the clusters leads to performance gains when the model is tested on tasks from the other clusters. Xu et al. (2022) continue this line of work by increasing the number of tasks used for pre-training from a few dozen to over 1000, showing that increasing the number of tasks is a good alternative to increasing the model size.
@@
Another possible approach for zero-shot classification is to use a large language model (LLM) to generate a synthetic dataset $\mathcal{S}$, tailored for the target task. Further, $\mathcal{S}$ can be used as training data in a fully-supervised setup, either to fine-tune an LLM or to train a conventional model from scratch. Although the dataset $\mathcal{S}$ can be made arbitrarily large, it can quickly hit diminishing returns, as it tends to contain redundant data Ye et al. (2022). Moreover, the performance obtained by using $\mathcal{S}$ is not as high as it would be when using a real human-labeled dataset of comparable size. To this end, several efforts have been made to improve the quality of $\mathcal{S}$ Meng et al. (2022); Ye et al. (2022). A generic dataset, such as Wikipedia, is sometimes employed as a source of additional diversity. Meng et al. (2022) use controlled text generation Hu et al. (2017) to direct an LLM to generate texts that are relevant for a specific class label. To increase sample diversity, repeating sequences are penalized. Alternatively, Yu et al. (2023) use Wikipedia as a general purpose corpus. In their method, called ReGen, a retrieval model is trained using contrastive learning for the task of finding the most relevant documents from the corpus. ProGen Ye et al. (2022) and ZeroGen Ye et al. (2022) represent other alternatives to generate synthetic datasets. ZeroGen uses prompt engineering to generate $\mathcal{S}$ from scratch, without the need of additional data. ProGen employs a feedback loop to iteratively refine a task-specific model.
@@ 5 Open-Set Text Classification and Discovery
Formal definitions. Open-set learning aims to classify samples from $C^{kk}$, while rejecting samples from $C^{uu}$ encountered during inference. The training set is defined a $\mathcal{D}=\{(x,y)\mid x\in\mathcal{X},y\in C^{kk}\}$, while the test set is given by $\mathcal{T}=\{(x,y)\mid x\in\mathcal{X},y\in C^{kk}\cup C^{uu}\}$. An open-set learning model is defined as $h:\mathcal{X}\to C^{kk}\cup\{\text{unknown}\}$, i.e. for any sample that belongs to $C^{uu}$, the model should label the respective sample as unknown.
In the open-set learning and discovery setup, the model is defined as $h:\mathcal{X}\to C^{kk}\cup C^{uu}$, i.e. the model must be able to classify any data sample, regardless of the fact that the sample belongs to $C^{kk}$ or $C^{uu}$. To classify samples into $C^{uu}$, classes belonging to $C^{uu}$ must be discovered during inference.
@@
There are a few methods that avoid estimating this threshold. Starting from the observation that there are spaces where feature vectors of instances from the same class generally reside close together, some researchers use methods specific to clustering. In this context, an outlierness factor can be calculated to decide if an instance should be classified into a class or allocated to the open space. If an example is an outlier, it probably belongs to the open space. Various outlierness factors have been proposed so far. Walkowiak et al. (2019a) employ the Local Outlier Factor (LOF), which uses a weighted Euclidean distance to find outliers based on the distance to their local neighbors. Subsequent studies of the same authors Walkowiak et al. (2019b, 2020) use LOF as well. Another factor, called Angle-Based Outlier Factor, which was originally introduced by Kriegel et al. (2008), is later adapted for text classification by Walkowiak et al. (2018), under the name ABOF2.
@@
Another solution to avoid setting an arbitrary threshold is to identify the examples belonging to open space before performing the classification task. The problem is therefore broken into two subtasks, outlier detection (OD) and closed-set classification, which can be treated independently. In this setup, the classification part is closed, hence OD is the difficult part.
@@
Open-set solutions via outlier / novelty detection. Kannan et al. (2017) divide the outlier detection algorithms for text in three subcategories: distance-based, density-based, and subspace-based. More recently, energy-based models have also been explored as a promising method for detecting out-of-distribution data. Grathwohl et al. (2019) propose a Joint Energy-Based Model (JEM) by reinterpreting a softmax-based classification network as a generative model. This approach uses the logits to define an energy-based representation of the joint distribution of data points and labels. Another notable contribution to outlier detection is the typicality test, introduced by Nalisnick et al. (2019). This method determines whether a given data point belongs to the model’s typical set, rather than relying solely on likelihood estimation. The test is based on the observation that deep generative models can assign higher likelihoods to OOD data than to in-distribution samples. By focusing on typicality instead of raw probability density, this method offers a more robust approach to detect outliers.
@@
We emphasize that there are two main types of distribution shifts Baran et al. (2023): semantic shift, defined as the occurrence of new classes, and background shift, defined as changes that are class-agnostic, such as the level of formality, stance, and so on. If the OD detects background shift, the examples belonging to $C^{kk}$ that exhibit background shift will not even make it to the classifier. Consequently, the outlier detector should only detect semantic shift, but this is difficult without access to outliers during training. The subject of semantic versus background shift in text is discussed in detail by Arora et al. (2021). While outlier detection in general is very well studied Wang et al. (2019), there are very few articles focused on text, especially ones trying to isolate semantic shift. Yet, a simple way to detect semantic shift is to use rare terms Mohotti and Nayak (2020).
@@
Chen et al. (2024) use adversarial disagreement maximization to increase the difference between in-distribution and OOD examples, improving on their previous solution Chen et al. (2023). Kim et al. (2022) propose a harder variant of OSTC, that is zero-shot. In this new setup, there are no labeled examples at all. The training data consists of a set of $n$ unlabeled texts $x_{1},...,x_{n}$ that must be classified into $k$ classes $c_{1},...,c_{k}$. Each class $c_{i}$ is only specified through a set of words. The proposed method obtains better results than the corresponding closed-set solutions Meng et al. (2020); Wang et al. (2021), when tested in an open-set setting.
@@
6 Continual Learning from Text
Formal definition. Continual learning (CL) is a paradigm in machine learning focused on enabling systems to continuously learn from a stream of data over time. Unlike traditional approaches that assume a static data distribution and rely on fixed datasets, CL addresses challenges such as retaining knowledge from previous tasks (avoiding catastrophic forgetting) and adapting to new tasks without extensive retraining Parisi et al. (2019). CL can be defined as the process of training a model on a sequence of tasks, where each task $t$ can be characterized by a distinct data distribution. At any point in time, the model receives a batch of training samples $D_{t,b}=\{(x,y)\mid x\in\mathcal{X}_{t,b},y\in\mathcal{Y}_{t,b}\}$, where $\mathcal{X}_{t,b}$ represents a subset of input samples and $\mathcal{Y}_{t,b}$ are the corresponding (task-specific) labels. The task identity is given by $t\in\{1,$$\infty\}$, while the batch index is represented by $b\in\{1,$$\infty\}.$ Theoretically, the continual learning process continues indefinitely during operation, hence the indexes $t$ and $b$ can grow endlessly. A task $t$ is formally defined by its data distribution $P(\mathcal{X}_{t},\mathcal{Y}_{t})$ Wang et al. (2024).
Trends and approaches. CL is a relevant framework for open-set learning, as it can address challenges related to the open-set scenarios, especially in tasks such as text classification by topic when the topic distribution changes. These challenges include handling unknown classes and adapting to an evolving data distribution. CL can help to discover and learn new topics during inference. Both Chen and Liu (2014) and Gupta et al. (2020) propose frameworks that enable models to accumulate knowledge over time, leveraging past knowledge to improve topic extraction, while mitigating catastrophic forgetting. These approaches integrate lifelong knowledge retention and transfer, ensuring that previously learned topics inform future learning tasks. By incorporating prior knowledge, these models enhance topic coherence, particularly in sparse data scenarios, where limited labeled data can affect the quality of extracted topics. Chen and Liu (2014) introduce the Lifelong Topic Model (LTM), which mines prior knowledge from past document collections and integrates it into new topic modeling tasks through probabilistic inference. Similarly, Gupta et al. (2020) propose the Lifelong Neural Topic Modeling (LNTM) framework, which relies on neural networks and selective data augmentation to balance stability and plasticity. LTM is designed for multi-domain topic extraction, identifying topic structures across different domains, while LNTM is more suitable for continuous document streams, maintaining topic stability, while adapting to novel information. Moreover, LTM employs a knowledge-based topic modeling approach that refines topic quality through extracted knowledge sets, whereas LNTM integrates topic regularization and selective data augmentation to enhance adaptation, while preventing excessive forgetting.
@@ 7 Conclusion and Future Work
Future directions. Continual learning is likely one of the most promising paths towards unlocking open-set learning and discovery in the text domain. However, adapting continual learning techniques to mitigate open-set learning and discovery requires addressing additional complexities, such as novelty detection, adaptive knowledge retention, and dynamic model scalability. One promising direction is disentanglement-based regularization, which separates learned representations into two spaces: a stable and generic one, and a flexible and task-specific one. This method enhances knowledge retention without restricting the adaptability of the model to new tasks, as demonstrated by Huang et al. (2021). Another important avenue is optimizing memory selection for replay, where approaches such as k-means clustering can be used to retain only the most representative examples from previous tasks.
@@
Research on continual learning for natural language generation (NLG) has also provided insights into how adaptive architectures can support evolving textual domains. Yang et al. (2022) propose a transformer calibration mechanism to minimize interference between tasks in continual learning scenarios for NLG. This technique leverages attention calibration and feature recalibration to enable language models to adapt incrementally to new tasks, while preserving previously learned knowledge. Similar strategies could be beneficial for open-set continual learning in text classification, where models must adjust their representation space dynamically as new topics emerge.
@@
Sun et al. (2020) introduce LAMOL (Language Modeling for Lifelong Language Learning), which reformulates various NLP tasks into language modeling. Under this framework, the model generates the replay samples from previously learned tasks, reducing catastrophic forgetting without storing large historical datasets. Although originally designed for a closed set of tasks, LAMOL could be extended to open-set scenarios by incorporating a novelty detection module. Whenever the system encounters anomalous samples or unfamiliar topics, it can trigger an incremental update process that treats the new categories or subject area as separate tasks. This setup would allow the model to enlarge its internal knowledge base, maintaining prior knowledge, while adapting dynamically to unseen class distributions.
@@
Drawing inspiration from CoLeCLIP Li et al. (2024), which introduces open-domain continual learning in the image domain through a combination of task-specific prompting and joint vocabulary learning, these principles can be adapted for text classification in open-set scenarios. The authors propose the development of shared representations and an adaptive prompting mechanism that enable a model to adjust to new tasks and classes as they emerge, while preserving previously acquired knowledge. Applied to text, this approach translates into designing dynamic prompts and an extensible vocabulary that guide classification under shifting class distributions, simultaneously detecting the emergence of new or unknown topics. For instance, building on the techniques employed by CoLeCLIP, a language model could learn task-specific prompts for different topics, continually adapting and expanding its internal vocabulary as it encounters new terms or concepts. Such an approach would likely need to involve a regularization mechanism to prevent catastrophic forgetting, employing selective memory strategies to retain critical information from past data.
8 Limitations

### v3

@@ 5 Zero-Shot Text Classification
When no examples from the target classes are available, various alternative tasks can be leveraged, with or without additional training, to perform zero-shot text classification. For instance, Yin et al. (2019) propose using the entailment task as a zero-shot text classification task. To decide if a text $x$ should be classified into a certain class, e.g. “politics”, one can ask if $x$ entails “The previous text is about politics”. This procedure is applied for each of the possible classes, and the one with the best confidence is considered the correct one. Using this technique, any model capable of solving the entailment problem can implicitly be used to perform zero-shot text classification. The zero-shot approach based on entailment (Yin et al., 2019) represents the basis for how large language models (LLMs) are used via prompting to solve various downstream tasks. LLMs (Bommasani et al., 2021; Yang et al., 2025; Zhou et al., 2024) are usually pre-trained on a very large corpus of unlabeled data, using a variety of self-supervised tasks, e.g. next token prediction, sentence order prediction, etc. During pre-training, LLMs learn general facts, such as language structure, making them ideal for task adaptation. LLMs are subsequently used in different downstream tasks, which often involve zero-shot setups. There are two main methodologies for adapting LLMs to a target task: fine-tuning and instruction tuning.
Fine-tuning involves adapting a model to a task by performing an extra training step using task-specific data. Instruction tuning, defined as fine-tuning language models on a collection of datasets described via instructions, can be seen as a generalization over the framework proposed by Yin et al. (2019). Instruction tuning can lead to better performance on unseen tasks, if the target task is described via simple instructions provided in natural language. Models trained with instruction tuning can perform various tasks, including zero-shot classification, as shown by Zhang et al. (2024). This capability is achieved by training language models to respond to simple natural language commands, with some degree of generality. For instance, Wei et al. (2022) and Sanh et al. (2022) explore similar ideas, dividing the tasks for which datasets are available into separate clusters, each containing multiple datasets. Fine-tuning via instructions on any of the clusters leads to performance gains when the model is tested on tasks from the other clusters. Xu et al. (2022) continue this line of work by increasing the number of tasks used for pre-training from a few dozen to over 1000, showing that increasing the number of tasks is a good alternative to increasing the model size.
@@
Another possible approach for zero-shot classification is to use an LLM to generate a synthetic dataset $\mathcal{S}$, tailored for the target task. Further, $\mathcal{S}$ can be used as training data in a fully-supervised setup, either to fine-tune an LLM or to train a conventional model from scratch. Although the dataset $\mathcal{S}$ can be made arbitrarily large, it can quickly hit diminishing returns, as it tends to contain redundant data (Ye et al., 2022). Moreover, the performance obtained by using $\mathcal{S}$ is not as high as it would be when using a real human-labeled dataset of comparable size. To this end, several efforts have been made to improve the quality of $\mathcal{S}$ (Meng et al., 2022; Ye et al., 2022). A generic dataset, such as Wikipedia, is sometimes employed as a source of additional diversity. Meng et al. (2022) use controlled text generation (Hu et al., 2017) to direct an LLM to generate texts that are relevant for a specific class label. To increase sample diversity, repeating sequences are penalized. Alternatively, Yu et al. (2023) use Wikipedia as a general purpose corpus. In their method, called ReGen, a retrieval model is trained using contrastive learning for the task of finding the most relevant documents from the corpus. ProGen (Ye et al., 2022) and ZeroGen (Ye et al., 2022) represent other alternatives to generate synthetic datasets. ZeroGen uses prompt engineering to generate $\mathcal{S}$ from scratch, without the need of additional data. ProGen employs a feedback loop to iteratively refine a task-specific model.
@@ 6 Open-Set Text Classification and Discovery
There are a few methods that avoid estimating this threshold. Starting from the observation that there are spaces where feature vectors of instances from the same class generally reside close together, some researchers use methods specific to clustering. In this context, an outlierness factor can be calculated to decide if an instance should be classified into a class or allocated to the open space. If an example is an outlier, it probably belongs to the open space. Various outlierness factors have been proposed so far. Walkowiak et al. (2019a) employ the Local Outlier Factor (LOF), which uses a weighted Euclidean distance to find outliers based on the distance to their local neighbors. Subsequent studies of the same authors (Walkowiak et al., 2020, 2019b) use LOF as well. Another factor, called Angle-Based Outlier Factor, which was originally introduced by Kriegel et al. (2008), is later adapted for text classification by Walkowiak et al. (2018), under the name ABOF2.
@@
Another solution to avoid setting an arbitrary threshold is to identify the examples belonging to the open space before performing the classification task. The problem is therefore broken into two subtasks, outlier detection (OD) and closed-set classification, which can be treated independently. In this setup, the classification part is closed, hence OD is the difficult part.
@@
Open-set solutions via outlier / novelty detection. Kannan et al. (2017) divide the outlier detection algorithms for text into three subcategories: distance-based, density-based, and subspace-based. More recently, energy-based models have also been explored as a promising method for detecting out-of-distribution data. Grathwohl et al. (2019) propose a Joint Energy-Based Model (JEM) by reinterpreting a softmax-based classification network as a generative model. This approach uses the logits to define an energy-based representation of the joint distribution of data points and labels. Another notable contribution to outlier detection is the typicality test, introduced by Nalisnick et al. (2019). This method determines whether a given data point belongs to the model’s typical set, rather than relying solely on likelihood estimation. The test is based on the observation that deep generative models can assign higher likelihoods to OOD data than to in-distribution samples. By focusing on typicality instead of raw probability density, this method offers a more robust approach to detecting outliers.
@@
We emphasize that there are two main types of distribution shifts (Baran et al., 2023): semantic shift, defined as the occurrence of new classes, and background shift, defined as changes that are class-agnostic, such as the level of formality, stance, and so on. If the OD detects background shift, the examples belonging to $C^{kk}$ that exhibit background shift will not even make it to the classifier. Consequently, the outlier detector should only detect semantic shift, but this is difficult without access to outliers during training. The subject of semantic vs. background shift in text is discussed in detail by Arora et al. (2021). While outlier detection in general is very well studied (Wang et al., 2019), there are very few articles focused on text, especially ones trying to isolate semantic shift. Yet, a simple way to detect semantic shift is to use rare terms (Mohotti and Nayak, 2020).
@@
Chen et al. (2024) use adversarial disagreement maximization to increase the difference between in-distribution and OOD examples, improving on their previous solution (Chen et al., 2023). Kim et al. (2022) propose a harder variant of OSTC, that is zero-shot. In this new setup, there are no labeled examples at all. The training data consists of a set of $n$ unlabeled texts $x_{1},...,x_{n}$ that must be classified into $k$ classes $c_{1},...,c_{k}$. Each class $c_{i}$ is only specified through a set of words. The proposed method obtains better results than the corresponding closed-set solutions (Meng et al., 2020; Wang et al., 2021), when tested in an open-set setting. LLMs have not been extensively studied for OSTC. Chen et al. (2024) are among the few to compare small vs. large language models on OOD detection and OSTC. They found that LLaMA2-7B performs poorly on both tasks, in both zero-shot and few-shot scenarios. Their study shows that prompting LLMs is not sufficient to accurately address OOD detection and OSTC, since a customized approach based on a smaller model can easily surpass zero-shot and few-shot LLM prompting.
@@
In the context of open-set text classification and discovery, instruction models can be used to perform the open-set text classification. However, it is difficult to discover new classes via prompting due to the large quantity of text and the relatively short-term memory of LLMs. However, using a reasoning model to interactively refine a list of possible topics is a path worth exploring.
7 Comparison of Learning Frameworks
In Table 2, we compare the three learning frameworks in terms of several key characteristics. From the comparative analysis, we next derive the main limitations of each learning framework.
Learning with background class. The performance depends heavily on the quality and diversity of background samples. Since it is impossible to sample from UUCs, the background data will never fully represent the real open-world. Hence, UUCs will likely be treated as one of the KKCs.
Zero-shot learning. Zero-shot models are sensitive to label wording, an aspect that is important to build representative class prototypes for UKCs. Without explicit supervision, the model might struggle to make subtle distinctions between fine-grained UKCs. UUC samples must map to one of the supplied classes, so zero-shot models cannot handle UUCs.
Open-set learning and discovery. The learning problem is hard. It is difficult to define boundaries around KKCs, without examples of UUCs. It is even more difficult to classify UUCs without labeled samples. Moreover, it is hard to solve open-set learning and discovery via a single objective, requiring multiple optimization stages.
Characteristic
Universum
Zero-shot
Open-set
Can detect OOD data
Set of classes is unbounded
Can classify OOD data
OOD training data required
Open-world assumption
Easy to optimize
Table 2: Comparison of learning frameworks in terms of various characteristics.
@@ 8 Conclusion and Future Work
Future directions. Continual learning is likely one of the most promising paths towards unlocking open-set learning and discovery in the text domain. However, adapting continual learning techniques to mitigate open-set learning and discovery requires addressing additional complexities, such as novelty detection, adaptive knowledge retention, and dynamic model scalability. Liu et al. (2023) identify several challenges in the development of an open-word continuous learning system, such as autonomous novelty detection, automatic acquisition of labeled data via interaction, and risk assessment during self-initiated adaptation. One promising direction is disentanglement-based regularization, which separates learned representations into two spaces: a stable and generic one, and a flexible and task-specific one. This method enhances knowledge retention without restricting the adaptability of the model to new tasks, as demonstrated by Huang et al. (2021). Another important avenue is optimizing memory selection for replay, where approaches such as k-means clustering can be used to retain only the most representative examples from previous tasks.
@@
Research on continual learning for natural language generation (NLG) has also provided insights into how adaptive architectures can support evolving textual domains. Yang et al. (2022) propose a transformer calibration mechanism to minimize interference between tasks in continual learning scenarios for NLG. This technique leverages attention calibration and feature recalibration to enable language models to incrementally adapt to new tasks, while preserving previously learned knowledge. Similar strategies could be beneficial for open-set continual learning in text classification, where models must adjust their representation space dynamically as new topics emerge.
@@
Sun et al. (2020) introduce LAMOL (Language Modeling for Lifelong Language Learning), which reformulates various NLP tasks into language modeling. Under this framework, the model generates the replay samples from previously learned tasks, reducing catastrophic forgetting (Kirkpatrick et al., 2017) without storing large historical datasets. Although originally designed for a closed set of tasks, LAMOL could be extended to open-set scenarios by incorporating a novelty detection module. Whenever the system encounters anomalous samples or unfamiliar topics, it can trigger an incremental update process that treats the new categories or subject area as separate tasks. This setup would allow the model to enlarge its internal knowledge base, maintaining prior knowledge, while adapting dynamically to unseen class distributions.
@@
Drawing inspiration from CoLeCLIP (Li et al., 2025b), which introduces open-domain continual learning in the image domain through a combination of task-specific prompting and joint vocabulary learning, we can apply similar principles for text classification in open-set scenarios. The authors propose the development of shared representations and an adaptive prompting mechanism, which jointly enable a model to adjust to new tasks and classes as they emerge, while preserving previously acquired knowledge. Applied to text, this approach translates into designing dynamic prompts and an extensible vocabulary that guide classification under shifting class distributions, simultaneously detecting the emergence of new or unknown topics. For instance, building on the techniques employed by CoLeCLIP, a language model could learn task-specific prompts for different topics, continually adapting and expanding its internal vocabulary as it encounters new terms or concepts. Such an approach would likely need to involve a regularization mechanism to prevent catastrophic forgetting, employing selective memory strategies to retain critical information from past data.
While current open-set methods exhibit difficulties in threshold selection for open-space delimitation, LLMs offer significant potential for addressing class distribution shift. For instance, LLMs can provide auxiliary confidence signals by generating natural language explanations for classification decisions. When an LLM shows high uncertainty or produces inconsistent explanations for a sample, this signals potential inclusion into an unknown class, complementing traditional outlier detection methods. LLMs can also generate semantically coherent labels for discovered clusters (Figure 3, step 4), ensuring consistency with existing classes, e.g. proposing “sports” rather than “tennis” when existing classes are “culture”, “politics” and “economy”.
Acknowledgments
This work was supported by a grant of the Ministry of Research, Innovation and Digitization, CCCDI - UEFISCDI, project number PN-IV-P6-6.3-SOL-2024-0090, within PNCDI IV. This research is also supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 351416.
Limitations